OpenStudy (anonymous):

For the function f(x) = –2(x + 3)2 − 1, identify the vertex, domain, and range.

OpenStudy (anonymous):

can someone please help?

OpenStudy (anonymous):

\[f(x)=–2(x + 3)^2 − 1 ?\]

OpenStudy (anonymous):

\[f(x)=-2(x+3)^2-1=-2x^2-12x-19\] the vertex: x=\[\frac{ -b }{ 2a }\]=\[\frac{ -12 }{ 2*2}=-3\] y=-1 vertex >> (-3,-1) domain: all Real Numbers, \[-\infty \:<x<\infty \] ,\[\left(-\infty \:,\:\infty \:\right)\] range:\[\:(-\infty \:,\:-1]\]